### v1

@@ 2.2 Delve into Information Flow of Visual Tokens
i) Analysis of Information Flow We begin by analyzing information inflow and outflow patterns across vision encoder layers. Specifically, for an attention map $\mathbf{A}\in\mathbb{R}^{N\times N}$, we define information inflow for each token as its primary source of information ($\mathrm{argmax}(\mathbf{A}[i,:])$) and outflow as the primary destination of its information ($\mathrm{argmax}(\mathbf{A}[:,j])$). Figure 2 (Left) reveals two distinct information flow patterns: CLS tokens consistently engage in global information exchange while providing information to patch tokens, whereas patch tokens exhibit short-range, local, uniform flows in shallow layers and long-range, global, deterministic flows in deeper layers.
@@ 3.3 Cumulative Flow Importance Tracking
SparseVLM (2024.10) 57.6 62.5 53.7 1721 83.6 68.9 75.6 56.1 50.5 55.8 31.5 66.1 96.6%
PDrop (CVPR2025) 57.3 62.9 56.8 1766 82.3 68.8 75.1 56.5 51.1 54.7 30.5 65.8 96.7%
VisionZip (CVPR2025) 59.3 63.0 57.3 1778 85.2 68.7 76.6 57.3 51.2 56.4 31.7 67.7 98.5%
@@
SparseVLM (2024.10) 56.0 60.0 51.1 1696 80.5 67.1 73.8 54.9 51.4 53.4 30.0 62.7 93.8%
PDrop (CVPR2025) 57.1 61.6 56.3 1664 82.3 68.3 72.9 56.6 51.0 53.3 30.8 61.9 95.1%
VisionZip (CVPR2025) 57.6 62.0 56.2 1759 83.2 68.9 75.6 56.8 51.6 54.9 32.1 64.8 97.2%
@@
SparseVLM (2024.10) 52.7 56.2 46.1 1505 75.1 62.2 68.2 51.8 50.1 51.1 23.3 57.5 84.6%
PDrop (CVPR2025) 47.5 58.8 50.5 1561 55.9 68.6 69.2 50.6 50.7 40.0 30.7 59.2 86.1%
VisionZip (CVPR2025) 55.1 60.1 55.3 1687 77.0 69.0 72.4 55.5 52.8 52.2 31.5 62.9 94.4%

### v2

@@ 2.2 Delve into Information Flow of Visual Tokens
i) Analysis of Information Flow
We begin by analyzing information inflow and outflow patterns across vision encoder layers. Specifically, for an attention map $\mathbf{A}\in\mathbb{R}^{N\times N}$, we define information inflow for each token as its primary source of information ($\mathrm{argmax}(\mathbf{A}[i,:])$) and outflow as the primary destination of its information ($\mathrm{argmax}(\mathbf{A}[:,j])$). Figure 2 (Left) reveals two distinct information flow patterns: CLS tokens consistently engage in global information exchange while providing information to patch tokens, whereas patch tokens exhibit short-range, local, uniform flows in shallow layers and long-range, global, deterministic flows in deeper layers.
@@ 3.3 Cumulative Flow Importance Tracking
SparseVLM (ICML25) 57.6 62.5 53.7 1721 83.6 68.9 75.6 56.1 50.5 55.8 31.5 66.1 96.6%
PDrop (CVPR25) 57.3 62.9 56.8 1766 82.3 68.8 75.1 56.5 51.1 54.7 30.5 65.8 96.7%
VisionZip (CVPR25) 59.3 63.0 57.3 1778 85.2 68.7 76.6 57.3 51.2 56.4 31.7 67.7 98.5%
@@
SparseVLM (ICML25) 56.0 60.0 51.1 1696 80.5 67.1 73.8 54.9 51.4 53.4 30.0 62.7 93.8%
PDrop (CVPR25) 57.1 61.6 56.3 1664 82.3 68.3 72.9 56.6 51.0 53.3 30.8 61.9 95.1%
VisionZip (CVPR25) 57.6 62.0 56.2 1759 83.2 68.9 75.6 56.8 51.6 54.9 32.1 64.8 97.2%
@@
SparseVLM (ICML25) 52.7 56.2 46.1 1505 75.1 62.2 68.2 51.8 50.1 51.1 23.3 57.5 84.6%
PDrop (CVPR25) 47.5 58.8 50.5 1561 55.9 68.6 69.2 50.6 50.7 40.0 30.7 59.2 86.1%
VisionZip (CVPR25) 55.1 60.1 55.3 1687 77.0 69.0 72.4 55.5 52.8 52.2 31.5 62.9 94.4%